OpenStudy (anonymous):

I am stuck on this question. find the slope-intercept equation of the line that has the given characteristics slope 9.5 and y-intercept (0,-5)

OpenStudy (anonymous):

Do you know about this equation y = mx + b

OpenStudy (anonymous):

We are going to use (y - y1) = m(x - x1) where y1 = -5; x1 = 0; m = 9.5 when then plug in the values into the equation above (y - -5) = 9.5(x - 0) y + 5 = 9.5x y = 9.5x - 5 That, is the slope intercept equation.